OpenStudy (littlebird):

Find the charge and stock name of each formula. Sn1O3 ( Sn1O2.5 ) Sn2O5 Sn1O2, stannic oxide ( Sn1O1.5 ) Sn2O3 Sn1O1 ( Sn1O0.5 ) Sn2O1

OpenStudy (cuanchi):

Sn can have +2 or +4 and O -2 oxidation state. the most common for Sn is +4 for the ions I don't know the names here are the charges SnO3 (+4) + 3(-2)= -2 (+2) + 3(-2) = -4 Sn2O5 (+4) 2 + 5 (-2) = -2 (+2) 2+ 5(-2) = -6 SnO2 (+4)+ 2 (-2) = 0 Tin (IV) oxide (+2) + 2(-2) = +4 Sn2O3 (+4) 2 + 3 (-2) = +2 (+2) 2+ 3(-2) = -2 SnO (+4) + (-2) = +2 (+2) + (-2) = 0 Tin (II) oxide Sn2O (+4) 2 + (-2) = +6 (+2) 2+ (-2) = +2